in youngs double slit experiment an interference pattern is obtained on a screen by a light of wavelength 6000 amstrong units coming from coherent sources S1 and S2. at certain point P on the screen, third dark fringe is formed. then the path difference S1P – S2P in microns is

Grade:12th pass

1 Answers

Agnibh Kumar
6 years ago
The answer will be 1.5 Microns.
 
S1P-S2P = (2n-1)λ/2 …... ( i )
 
here, λ = 6000 A and n = 3
 
Therefore, put the values in ( i ),
 
(2x3-1)6000/2 = 5x6000/2 = 15000 A = 1.5 microns